Topic: Engineering Mathematics
The probability that a teacher will give an unannounced test during any class is 1/5. If a student is absent twice, then probability that misses atleast one test is
- A. 2/3
- B. 4/5
- C. 7/25
- D. 9/25
Answer: D. 9/25
Explanation: The probability of a test on a given day is 1/5, so the probability of no test is 4/5. The probability that the student misses no test over two absent days is $(4/5) \times (4/5) = 16/25$. Therefore, the probability of missing at least one test is $1 - P(\text{miss no test}) = 1 - 16/25 = 9/25$.